OpenStudy (anonymous):

Solve. 2x^2 + x = 15.

OpenStudy (jlastino):

subtract 15 from both sides. You get a quadratic equation. Try to solve it using factoring, if all else fails, use the quadratic equation

OpenStudy (anonymous):

-3, 5/2

sam (.sam.):

2x^(2)+x=15 2x^(2)+x-15=0 (x+3)(2x-5)=0 x=-3,(5)/(2)